# Explainability of Large Language Models using SMILE: Statistical Model-agnostic Interpretability with Local Explanations

### v1

@@ 5.2.9 Token-Level Attention for Different Sentence Structures
Figure 12: Box plot showing attention scores for the tokens “meaning” and “life” across ten prompts with varying sentence structures. Hovering over each point reveals the full sentence associated with that score.
@@ 6 Limitations
While SMILE provides a flexible and model-agnostic framework for interpreting large language models (LLMs), its application introduces several limitations due to the use of perturbation-based sampling and LLMS’ black-box nature.
1) Perturbing prompts at the token level can lead to disproportionately large changes in LLM outputs, even when the semantic meaning remains largely intact. This sensitivity challenges the assumption of local smoothness required for regression-based attribution and can reduce explanation stability.
@@ 7 Conclusion
This work introduced a novel interpretability framework tailored for instruction-based text generation models. Leveraging SMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output. Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
Whereas most existing interpretability methods highlight important tokens in the model’s output [13, 134], SMILE shifts attention to the input space. By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
Our experiments show that SMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma. The resulting heatmaps and attribution scores intuitively understand model behaviour, empowering users to control better and refine prompt design.
This work lays the groundwork for more transparent and trustworthy natural language systems. Future directions include extending SMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
@@ 8 Future Works
The datasets used in this article are publicly available, and the code supporting this paper is published online on GitHub. For the SMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy. For the proposed algorithm introduced in this paper, see Sara068/CELL_SMILE.

### v2

@@ 5.2.8 Token-Level Attention for Different Sentence Structures
Figure 12: Box plot showing attention scores for the tokens “meaning” and “life” across ten prompts with varying sentence structures. Each point represents the attention score assigned to the token in a specific sentence instance.
@@ 6 Limitations
While gSMILE provides a flexible and model-agnostic framework for interpreting large language models (LLMs), its application introduces several limitations due to the use of perturbation-based sampling and LLMS’ black-box nature.
1) Perturbing prompts at the token level can lead to disproportionately large changes in LLM outputs, particularly in English, even when the semantic meaning remains largely intact. This sensitivity challenges the assumption of local smoothness required for regression-based attribution and can reduce explanation stability.
@@ 7 Conclusion
This work introduced a novel interpretability framework tailored for instruction-based text generation models. Leveraging gSMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output. Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
Whereas most existing interpretability methods highlight important tokens in the model’s output [63, 64], gSMILE shifts attention to the input space. By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma. The resulting heatmaps and attribution scores provide an intuitive understanding of model behavior, empowering users to control and refine prompt design better. gSMILE was systematically evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including fidelity, stability, and consistency.
This work lays the groundwork for more transparent and trustworthy natural language systems. Future directions include extending gSMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
@@ 8 Future Works
  6. 6.
    Optimisation of Computational Efficiency: To address the computational demands of gSMILE, future work will investigate dimensionality reduction techniques and approximation strategies. These approaches aim to reduce execution time while preserving the integrity and interpretability of the model’s outputs.
@@
The datasets used in this article are publicly available, and the code supporting this paper is published online on GitHub. For the gSMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy. For the proposed algorithm introduced in this paper, see Sara068/CELL_SMILE.